eclipsedstar (eclipsedstar):

Verify the given identity using double-angle formulas. Did I do this right?

OpenStudy (just_one_last_goodbye):

You got this one @AlexandervonHumboldt2 ? :)

OpenStudy (alexandervonhumboldt2):

you got wut to do JOLG?

OpenStudy (just_one_last_goodbye):

Nah bruh ^_^ I'll go help someone else. ^_^ good luck

OpenStudy (triciaal):

can't see the question.

eclipsedstar (eclipsedstar):

question: \[\frac{ cosx+\cos3x }{ 2\cos2x }=cosx\] work: \[\frac{ cosx+\cos(2x+x)}{ 2\cos2x }\] = \[\frac{ cosx+cos2xcosx-\sin^2x)-\sin^2xsinx}{ 2cos2x }\] = \[\frac{ cosx+(1-2\sin^2x)cosx-(2sinxcosx)sinx }{ 2\cos2x }\] =\[\frac{ cosx+cosx(1-2\sin) }{ 2\cos2x }\] =\[\frac{ 2cosx-4\sin^2xcosx }{ 2(1-2\sin^2x)}\] = cosx

eclipsedstar (eclipsedstar):

The page refreshed on me ;-;

eclipsedstar (eclipsedstar):

also I think I put in some typos...

OpenStudy (triciaal):